# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 2 Learning Multiclass Linear Classifiers with Adversarial Label Noise
The remaining part of the section is structured as follows: In Section 2.1, we state and prove the key structural result about the class of multiclass linear classifiers, Proposition 2.4, which will be exploited by our algorithm. In Section 2.2, we provide the analysis of Algorithm 2 with Proposition 2.8. In Section 2.3, we analyze Algorithm 1, providing the full proof of Theorem 2.1. Finally, in Section 2.4, we provide the proof of Lemma 2.19, a supporting lemma that is used in the proof of Proposition 2.4. The main result of this section is Theorem 2.1.
@@ 2.1 Structural Result for Multiclass Linear Classification
In this section, we prove our main structural result, Proposition 2.4, which enables us to learn multiclass linear classifiers efficiently under adversarial noise. In particular, we show that for a multiclass linear classifier $f$, if $f$ is far from any constant function by more that a constant multiple of the error of $f$ with respect to the noisy label $y$, then the first moment of at least one of the noisy label regions has a positive correlation with a vector in the subspace of weights.
We prove this by finding carefully chosen vectors $\mathbf{v}^{(i)}{\in}\mathbb{R}^{d}$, $i{\in}[K]$, such that $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(y=i)]$ is guaranteed to be positive. Note that this sum is equal to $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(f(\mathbf{x})=i)]$ plus an error term for instances when $y$ is not equal to $f(\mathbf{x})$. By carefully selecting these vectors, we can ensure that a small probability of error cannot corrupt the sum by much. Formally, we show the following:
@@
Define $\mathrm{OPT}\coloneqq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. First note that without loss of generality we can assume that $\mathbf{w}^{(1)}=\mathbf{0},t_{1}=0$. This can be seen by noting that the output label does not change if we subtract the same function from all affine functions that define the multiclass linear classifier. Specifically, by subtracting $\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$, where $j=\operatorname*{argmax}_{i\in[K]}t_{i}$ from all affine functions and renaming the labels, we ensure that one of the affine functions becomes $\mathbf{0}$ and that $t_{i}\leq 0$ for all $i\in[K]$.
We define $L_{i}:\mathbb{R}^{d}\to[K],L_{i}(\mathbf{x})=\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}$ the corresponding affine functions such that $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}L_{i}(\mathbf{x})$ and we define the function $s(\mathbf{x})\coloneqq\max_{i\in[K]}L_{i}(\mathbf{x})$.
For some parameter $\gamma>0$ which we will assign later, let $\mathbf{v}^{(i)}\coloneqq\mathbf{w}^{(i)}/(\gamma-t_{i})$ (this is well-defined as $\gamma>0$ and $t_{i}\leq 0$) for all $i\in[K]$. Note that for all nonzero vectors $\mathbf{v}^{(i)}$, we have that $L_{i}(\mathbf{x})\geq\gamma$ if and only if $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$. We can decompose the first moment of label $i$ as follows:
@@
where $S_{i}(\mathbf{x},y)=\mathds{1}(y\neq i,f(\mathbf{x})=i)$ and $T_{i}(\mathbf{x},y)=\mathds{1}(y=i,f(\mathbf{x})\neq i)$ for $i\in[K]$. Taking the inner product with $\mathbf{v}^{(i)}$ and summing Equation 1 for all $i\in[K]$, we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(y)}\cdot\mathbf{x}]$ $=\underbrace{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}]}_{I_{1}}\underbrace{\;-\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}S_{i}(\mathbf{x},y)]+\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(x,y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}T_{i}(\mathbf{x},y)]}_{I_{2}}\;.$
Note that $I_{1}$ is the first moment of $\mathbf{v}^{(f(\mathbf{x}))}$, i.e., the first moment of the noiseless multiclass linear classifier, and $I_{2}$ is the contribution that corresponds to the noise. Let $g(\mathbf{x})\coloneqq\max_{j\in[K]}\mathbf{v}^{(j)}\cdot\mathbf{x}$. We first bound $I_{1}$ from below.
@@
Note that it holds $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 0$ because $L_{f(\mathbf{x})}(\mathbf{x})=\max_{i}L_{i}(\mathbf{x})\geq L_{1}(\mathbf{x})=0$ and therefore $\mathbf{w}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq|t_{i}|$ which gives that $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 0$. Hence, by applying Markov’s inequality, we have
@@
where the last equality holds by the definition of $\mathbf{v}^{(i)}$. Note that the event $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$ is a superset of the event $g(\mathbf{x})\geq 1$. Indeed, if $g(\mathbf{x})\geq 1$, there exists some $i\in[K]$ such that $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$, which implies that $L_{i}(\mathbf{x})\geq\gamma$, and therefore $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$. This gives that $I_{1}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$ and completes the proof of Claim 2.5. ∎
@@
There exists a $\gamma>0$ such that $I_{2}\geq-C\left(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]}+\mathrm{OPT}\right)$, where $C>0$ is a universal constant, and $961/48\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]/$ 96.
@@
We show how to bound the contribution of the terms containing $S_{i}$, as similar arguments can be used to bound the contribution of the terms involving $T_{i}$. In order to bound the contribution of the $S_{i}$ terms from above, notice that
@@
$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\right]\;,$
where $S\coloneqq\sum_{i=1}^{K}S_{i}$. From the assumptions of Proposition 2.4, we have that $f$ has a missclassification error at most $\mathrm{OPT}$, and therefore we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$. Furthermore, only one of $S_{i}$ and $T_{i}$ for all $i\in[K]$ can be non-zero for each point $(\mathbf{x},y)$, which gives that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\leq 1$. Let $R(\mathbf{x})\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$ and note that from the above we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$ and $R(\mathbf{x})\in[0,1]$.
By the definition of the $\mathbf{v}^{(i)}$’s, we have $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\max_{i\in[K]}L_{i}(\mathbf{x})\geq\gamma]$. Note that $g(\mathbf{x})\geq 0$ by construction, therefore for any $t>0$, we have that
@@
$=\sum_{n=\lfloor t\rfloor}^{\infty}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]\;,$
where in the first equality, we exchange the order of the integrals using Tonelli’s theorem and in the inequality, we used that $\mathds{1}(g(\mathbf{x})\geq y)$ is a decreasing function of $y$. In order to bound the terms $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]$, we prove that the tails of $g$ are at least quadratically decreasing. We show the following:
@@
For the proof of Lemma 2.7, we refer the reader to Section 2.4. Now, we leverage Lemma 2.7 to complete the proof. Note that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma].$ Furthermore, since $t_{i}\leq 0$ for all $i\in[K]$, it follows that if $\mathbf{w}^{(i)}\geq n(\gamma-t_{i})$, then $\mathbf{w}^{(i)}\geq n\gamma-t_{i}$. This, in turn, implies that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq n]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma].$ Therefore, using Lemma 2.7, we have that
@@
where in the last inequality, we used the approximation of the sum by an integral. Therefore, by choosing $t=2+(\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]/\mathrm{OPT})^{1/2}$, we have that
@@
where $D^{\prime}$ is a reflection of distribution $D$ with respect to $\mathbf{0}$. As a result, we can arrive at the same bound as for the term defined by the $S_{i}$’s. Therefore, we conclude that the total contribution of outliers is at most $O(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]}+\mathrm{OPT})$, which completes the proof of Claim 2.6. ∎
By the choice of $\gamma$ in Claim 2.6 and the assumption that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq i]\geq C\cdot\mathrm{OPT}+\epsilon$ we have that $\mathrm{OPT}\leq({96}/C)\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$. Therefore, for a sufficiently large constant $C>0$, we get that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(y)}\cdot\mathbf{x}]\gtrsim\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$ and this gives us

### v2

@@ 2 Learning Multiclass Linear Classifiers with Adversarial Label Noise
The body of this section is structured as follows: In Section 2.1, we state and prove the key structural result about the class of multiclass linear classifiers, Proposition 2.4, which will be exploited by our algorithm. In Section 2.2, we provide the analysis of Algorithm 2 in Proposition 2.8. In Section 2.3, we analyze Algorithm 1, providing the full proof of Theorem 2.1. Finally, in Section 2.4, we provide the proof of Claim 2.19, a supporting lemma that is used in the proof of Proposition 2.4.
$\mathbf{v}^{(1)}$
$\mathbf{v}^{(2)}$$\mathbf{v}^{(1)}$
Figure 1: Illustration of two consecutive steps of our iterative subspace approximation procedure for 4-class Multiclass Linear Classification. In the left figure, the space $\mathbb{R}^{d}$ is partitioned according to one already discovered direction. The highlighted rectangular regions correspond to areas where a constant function over the current subspace $V$ yields an insufficient approximation, indicating that additional relevant directions can be found. In contrast, the non-highlighted regions already permit a sufficiently good constant approximation. The right figure illustrates that we discovered the second hidden direction so that partitioning the space to sufficiently small cubes leads to a good approximation.
@@ 2.1 Structural Result for Multiclass Linear Classification
In this section, we prove our main structural result, Proposition 2.4, which enables us to efficiently learn multiclass linear classifiers under adversarial noise. In particular, we show that for a multiclass linear classifier $f$, if $f$ is far from any constant function by more than a constant multiple of its error with respect to the noisy label $y$, then the first moment of at least one of the noisy label regions has a non-trivial correlation with a vector in the subspace of weights defining $f$.
We prove this by finding carefully chosen vectors $\mathbf{v}^{(i)}{\in}\mathbb{R}^{d}$, $i{\in}[K]$, such that $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(y=i)]$ is guaranteed to be non-trivially large. Note that this sum is equal to $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(f(\mathbf{x})=i)]$ plus an error term for instances when $y$ is not equal to $f(\mathbf{x})$. By carefully selecting these vectors, we ensure that even a small error probability does not significantly reduce the overall sum. Formally, we show the following:
@@
Define $\mathrm{OPT}\coloneqq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. Without loss of generality, we may assume that $\mathbf{w}^{(1)}=\mathbf{0}$ and $t_{1}=0$. This follows from the observation that subtracting the same function from every classifier does not change the output label. In particular, if we subtract $\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$, where $j\in\operatorname*{argmax}_{i\in[K]}t_{i}$, from all affine functions and then rename the labels accordingly, one of the resulting affine functions becomes identically $\mathbf{0}$ and we obtain $t_{i}\leq 0$ for all $i\in[K]$.
We define $L_{i}:\mathbb{R}^{d}\to\mathbb{R},L_{i}(\mathbf{x})=\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}$ the corresponding affine functions such that $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}L_{i}(\mathbf{x})$ and we define the function $s(\mathbf{x})\coloneqq\max_{i\in[K]}L_{i}(\mathbf{x})$.
For some parameter $\gamma>0$ which we will chose appropriately below, let $\mathbf{v}^{(i)}\coloneqq\mathbf{w}^{(i)}/(\gamma-t_{i})$ (this is well-defined as $\gamma>0$ and $t_{i}\leq 0$) for all $i\in[K]$. Let $g(\mathbf{x})\coloneqq\max_{i\in[K]}\mathbf{v}^{(i)}\cdot\mathbf{x}$. Note that for all nonzero vectors $\mathbf{v}^{(i)}$, we have that $L_{i}(\mathbf{x})\geq\gamma$ if and only if $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$ which means that $g(\mathbf{x})\geq 1$. Furthermore, because $\mathbf{w}^{(1)}=\mathbf{0}$ and $t_{1}=0$, we have that $g(\mathbf{x})\geq 0$ for all $\mathbf{x}\in\mathbb{R}^{d}$. We can decompose the first moment of label $i$ as follows:
@@
where $S_{i}(\mathbf{x},y)=\mathds{1}(y\neq i,f(\mathbf{x})=i)$ and $T_{i}(\mathbf{x},y)=\mathds{1}(y=i,f(\mathbf{x})\neq i)$ for $i\in[K]$. Taking the inner product of $\operatorname*{\mathbf{E}}[x\mathds{1}(y=i)]$ with $\mathbf{v}^{(i)}$ and summing Equation 1 over $i\in[K]$, we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(y)}\cdot\mathbf{x}]$ $=\underbrace{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}]}_{I_{1}}\underbrace{\;-\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}S_{i}(\mathbf{x},y)]+\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(x,y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}T_{i}(\mathbf{x},y)]}_{I_{2}}\;.$ (2)
Note that $I_{1}$ is the the contribution of each $\mathbf{v}^{(i)}$ over $\mathbf{x}$ in the region where $f(\mathbf{x})=i$, and $I_{2}$ is the contribution that corresponds to the noise. We first bound $I_{1}$ from below.
@@
Note that the random variable $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}$ is non-negative since $L_{f(\mathbf{x})}(\mathbf{x})=\max_{i}L_{i}(\mathbf{x})\geq L_{1}(\mathbf{x})=0$ which implies that $\mathbf{w}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq|t_{i}|$ and consequently, $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 0$. Hence, by applying Markov’s inequality, we have
@@
where the last equality holds by the definition of $\mathbf{v}^{(i)}$. It remains to show that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$. To see this, note that the event $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$ contains the event $g(\mathbf{x})\geq 1$. If $g(\mathbf{x})\geq 1$ then by definition there exists some $i\in[K]$ such that $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$, which implies that $L_{i}(\mathbf{x})\geq\gamma$ (recall that $\mathbf{v}^{(i)}=\mathbf{w}^{(i)}/(\gamma-t_{i})$), and consequently, $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$. This gives that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$ and completes the proof of Claim 2.5. ∎
@@
There exists an absolute constant $C^{\prime}>0$ and $\gamma>0$ such that
$I_{2}\geq-C^{\prime}\left(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]}+\mathrm{OPT}\right)\;,$
and $961/48\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]/$ 96.
@@
We show how to bound the contribution of the terms involving $S_{i}$; similar arguments can be used to bound the contribution of the terms involving $T_{i}$. In order to bound the contribution of the $S_{i}$ terms we can write
@@
$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\right]$
$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[g(\mathbf{x})\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\right]\;,$
where $S=\sum_{i=1}^{K}S_{i}$. Since $f$ has misclassification error $\mathrm{OPT}$, we have $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$. Moreover, for each point $(\mathbf{x},y)$, at most one among $\{S_{i},T_{i}:i\in[K]\}$ can be nonzero which implies that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\leq 1$. Define $R(\mathbf{x})\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$. Then, by the law of total expectation, we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$ and $R(\mathbf{x})\in[0,1]$.
Recall that $g(\mathbf{x})\geq 0$ by construction, hence for any $t>0$,
@@
$=\sum_{n=\lfloor t\rfloor}^{\infty}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]\;,$ (3)
where in the first equality, we exchange the order of the integrals using Tonelli’s theorem and in the inequality, we used that $\mathds{1}(g(\mathbf{x})\geq y)$ is a decreasing function of $y$. In order to bound the terms $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]$, we show that the tails of $g$ decay at least quadratically. We show the following:
@@
For the proof of Lemma 2.7, we refer the reader to Section 2.4. Now, we leverage Lemma 2.7 to complete the proof. Recall that the function $s(\mathbf{x})$ of Lemma 2.7 is defined as $s(\mathbf{x})=\max_{i}L_{i}(\mathbf{x})$. We need to show that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq n]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]$. First, note that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma],$ by construction. Furthermore, recall that $g(\mathbf{x})=\max_{i\in[K]}\mathbf{v}^{(i)}\cdot\mathbf{x}$ where $\mathbf{v}^{(i)}=\mathbf{w}^{(i)}/(\gamma-t_{i})$ and that $L_{i}(\mathbf{x})=\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}$. Since $t_{i}\leq 0$ for all $i\in[K]$, it follows that if $\mathbf{w}^{(i)}\cdot\mathbf{x}\geq n(\gamma-t_{i})$, then $\mathbf{w}^{(i)}\cdot\mathbf{x}\geq n\gamma-t_{i}$. This, in turn, implies that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq n]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma].$ Therefore, using Lemma 2.7 into Equation 3, we have that
@@
where in the last inequality we used the approximation of the sum by an integral. Therefore, by choosing $t=2+(\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]/\mathrm{OPT})^{1/2}$, we have that
@@
where $D^{\prime}$ is a reflection of distribution $D$ with respect to $\mathbf{0}$. As a result, we can arrive at the same bound as for the term defined by the $S_{i}$’s. Therefore, we conclude that the total contribution of the noise is at most $O(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]}+\mathrm{OPT})$, which completes the proof of Claim 2.6. ∎
Now we show that $I_{2}\geq-\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]/2$. Recall that by the assumptions of Proposition 2.4 we have
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}f(\mathbf{x})\neq i\big{]}\geq C\,\mathrm{OPT}+\epsilon\quad\text{for all }i\in[K].$
We show that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\neq 0\big{]}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}f(\mathbf{x})\neq 1\big{]}$. To see this, observe that for any $i\in[K]$ for which $L_{i}$ is not identically zero, we have that $L_{i}(\mathbf{x})\neq 0$ almost surely (with respect to the Gaussian measure). Moreover, if for some $i\in[K]$ the function $L_{i}$ is identically zero and $i\neq 1$, then without loss of generality we may assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i]=0$, since we could set $1$ to be the label with minimum index in the set $\operatorname*{argmax}_{i\in[K]}t_{i}$ (for the original biases $t_{i}$). Hence, in every case for any $i\neq 1$ we have $L_{i}(\mathbf{x})\neq 0$ almost surely. Therefore, if $f(\mathbf{x})\neq 1$ then, except for a measure zero set it holds that $L_{i}(\mathbf{x})>0$ for some $i\neq 1$, which implies that $s(\mathbf{x})\neq 0$. As a result, it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\neq 0\big{]}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}f(\mathbf{x})\neq 1\big{]}$. Hence, it follows that
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\neq 0\big{]}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}f(\mathbf{x})\neq 1\big{]}\geq C\,\mathrm{OPT}+\epsilon.$
By the choice of $\gamma$ in Claim 2.6, we also have
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\geq\gamma\big{]}\geq\frac{1}{96}\,\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\neq 0\big{]}.$
Since $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}g(\mathbf{x})\geq 1\big{]}=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\geq\gamma\big{]}$, we get that that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\geq({C\,\mathrm{OPT}+\epsilon})/{96}$. Finally, by choosing $C$ sufficiently large (in particular, ensuring that $C\geq 256/((C^{\prime})^{2}+1)$), from Claim 2.6, we conclude that
$I_{2}\geq-\frac{1}{2}\,\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}g(\mathbf{x})\geq 1\big{]}\;.$
and that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}g(\mathbf{x})\geq 1\big{]}\gtrsim\epsilon$. Combining the above with Equation 2 and Claim 2.5, we obtain